Capacitors of 2μF and 6μF are connected to a 100V battery. How much charge I supplied by the battery in charging the capacitors when (a) they are connected in series and (b) when they are connected in parallel?


1
Expert's answer
2022-03-20T18:49:45-0400

(a)

"C_{net}=\frac{C_1C_2}{C_1+C_2}=\frac{2*6}{2+6}=1.5\:\rm \mu F"

"q_{net}=C_{net}V=1.5*10^{-6}*100=1.5*10^{-4}\:\rm C"

(b)

"C_{net}=C_1+C_2=2+6=8\:\rm \mu F"

"q_{net}=C_{net}V=8*10^{-6}*100=8*10^{-4}\:\rm C"